Kingbird Company provides you with the following condensed balance sheet information Assets Current assets Equity investments Equipment (net) Intangibles $ 43,900 64,800 238,100 58,300 $405,100 Total assets Liabilities and Stockholders Equity Current and long-term liabilities Stockholders equity $102,000 Common stock ($5 par) Paid-in capital in excess of par Retained earnings $ 19,900 112,800 170,400 303,100 Total liabilities and stockholders equity $405,100 For each transaction below, indicate the dollar impact (if any) on the following five items: (1) total assets, (2) common stock, (3) paid-in capital in excess of par, (4) retained earnings, and (5) stockholders equity. (Each situation is independent.) (a) Kingbird declares and pays a $0.60 per share cash dividend (1) Total assets (2) Common stock (3) Paid-in capital in excess of par (4) Retained earnings (5) Total stockholders equity(b) Kingbird declares and issues a 10% stock dividend when the market price of the stock is $15 per share (1) Total assets (2) Common stock (3) Paid-in capital in excess of par (4) Retained earnings (5) Total stockholders equity (c) Kingbird declares and issues a 30% stock dividend when the market price of the stock is $14 per share (1) Total assets (2) Common stock (3) Paid-in capital in excess of par (4) Retained earnings (5) Total stockholders equity (d) Kingbird declares and distributes a property dividend. Kingbird gives one share of its equity investment (ABC stock) for every two shares of Kingbird Company stock held. Kingbird owns 10,800 shares of ABC. ABC is selling for $9 per share on the date the property dividend is declared (1) Total assets (2) Common stock (3) Paid-in capital in excess of par (4) Retained earnings (5) Total stockholders equity(e) Kingbird declares a 3-for-1 stock split and issues new shares (1) Total assets (2) Commo (3) Paid-in capital in excess of par (4) Retained earnings (5) Total stockholders equity
Answer #1

Ans-a- Kingbird declares and pays a $0.60 per share cash dividend:-

Cash Dividend=$19,900/5 *0.60= $2,388

1 Total assets decrease 2,388
2 Common stock no effect 0
3 Paid-in capital in excess of par no effect 0
4 Retained earnings decrease 2,388
5 Total stockholders' equity decrease 2,388

b-Kingbird declares and issues a 10% stock dividend when the market price of the stock is $15 per share.

No. of shares outstanding = $19,900/5= 3,980

Stock dividend in share= 3,980*10%= 398

Stock dividend= 398* $15= 5,970

Common Stock Increase= 398*5= 1,990

Paid-in Capital in excess of par increase= 398* (15-5)= 3,980

Retained earnings decrease= 5,970

1 Total assets no effect 0
2 Common stock increase 1,990
3 Paid-in capital in excess of par increase 3,980
4 Retained earnings decrease 5,970
5 Total stockholders' equity no effect 0

c-Kingbird declares and issues a 30% stock dividend when the market price of the stock is $14 per share:-

No. of shares outstanding= $19,900/5= 3,980

Stock dividend in share= 3,980*30%= 1,194

Since it is large stock dividend as it is greater than 25%

Stock dividend in dollar= 1,194*5= 5,970

Common stock increase= 1,194*5= 5,970

Retained earnings decrease= 5,970

1 Total assets no effect 0
2 Common stock increase 5,970
3 Paid-in capital in excess of par no effect 0
4 Retained earnings decrease 5,970
5 Total stockholders' equity no effect 0

d- Kingbird declares and distributes a property dividend. Kingbird gives one share of its equity investment (ABC stock) for every two shares of Kingbird Company stock held. ABC is selling for $9 per share on the date the property dividend is declared.

No. of share outstanding = 19,900/5 = 3,980

Cost per share= 64,800/10,800= $6 per share

Dividend= 3,980/2*6= 11,940

1 Total assets decrease 11,940
2 Common stock no effect 0
3 Paid-in capital in excess of par no effect 0
4 Retained earnings decrease 11,940
5 Total stockholders' equity decrease 11,940

e- Kingbird declares a 3- for -1 stock split and issue new shares in stock split:-

No.of share outstanding = 19,900/5= 3,980

No.of new stock outstanding in share= 3,980*3= 11,940

There is no change in value since it will reduce the par value of share from $5 to 5/3

1 Total assets no effect 0
2 Common stock no effect 0
3 Paid-in capital in excess of par no effect 0
4 Retained earnings no effect 0
5 Total stockholders' equity no effect 0